the area of to the nearest hundredth of a 10° sector in a circle with diameter 12 centimeters.
~thank you so much

Answer by stanbon(75887)

the area of (???) to the nearest hundredth of a 10° sector in a circle with diameter 12 centimeters.
------------------------
Area of the circle = (pi)6^2 = 36pi
---
A 10 decree sector has 10/360 of the area
---
(10/360)(36pi) = (1/36)(36pi) = pi sq. cm.
====================================================
Cheers,
Stan H.
===========